OpenStudy (anonymous):

how to integrate 12/(2x+1)*2 not by substitution.

OpenStudy (anonymous):

first take the constant out \[\int\limits_{?}^{?}\frac{ 12 }{ (2x+1)^2 }dx=12\int\limits_{?}^{?}\frac{ 1 }{ (2x+1)^2 }dx\]

OpenStudy (anonymous):

you there?

OpenStudy (anonymous):

yeah thanks..:)

OpenStudy (anonymous):

do you know how to continue?

OpenStudy (anonymous):

yeah i do. :)

OpenStudy (anonymous):

I got the result already if you want us to compare

OpenStudy (anonymous):

it's -6/2x+1?

OpenStudy (anonymous):

oh and +7

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

thanks :)

OpenStudy (anonymous):

-6/2x+1 i dont know about the +7

OpenStudy (anonymous):

\[12 \int\limits_{}^{} (2x+1)^-2 dx = \frac{ 1 }{ 2 } \frac{ (2x+1}^-2{ -1 } = \frac{ -6 }{ 2x+1 }+c\]